OpenStudy (anonymous):

I've never done exponential notation before and its completely boggling my mind, how would you go about solving (81x^-4y^8)^-1/4(xy^1/4) ?

OpenStudy (anonymous):

can you use equation button down there?

OpenStudy (anonymous):

\[(81x ^{-4}-4y ^{8})^{-1/4}(xy ^{1/4})\]

OpenStudy (anonymous):

OK, you mean simplify not solve. Do you know the rules for powers?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

It's more straightforward if you know them, want me to explain them?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

First one \[a^p * a^q = a^{p+q}\]

OpenStudy (anonymous):

its pretty much a foreign language. im terrible at math

OpenStudy (anonymous):

How about 2^2 + 2^3 = 2^(2+3) = 2^5 (concrete example of Rule1

OpenStudy (anonymous):

okay that makes sense

OpenStudy (anonymous):

It shouldn't cos I accidentally put a plus instead of a *, lol. So you do a ^2 * b^5 = ?

OpenStudy (anonymous):

Do you want me to find you a site where u can study the rules? (8 of them).

OpenStudy (anonymous):

im taking a placement test to enroll in courses for college this fall

OpenStudy (anonymous):

Once u know the rules, powers are pretty easy.